OpenStudy (anonymous):

Find the midpoint of segment AG: A(12, 3) and G(8, 8). (4, -5) (20, 11) (2.5, 10) (10, 5.5)

OpenStudy (anonymous):

(10, 5.5) is the answer

OpenStudy (anonymous):

how is this? take the x's and divide by 2, and take the y's and divide by 2

OpenStudy (anonymous):

Using vectors \[AG= \left(\begin{matrix}8-12 \\ 8-3\end{matrix}\right)= \left(\begin{matrix}-4 \\ 5\end{matrix}\right)\] M mid point \[AM= \left(\begin{matrix}-4 \\ 5\end{matrix}\right)/2\] so \[M \left(\begin{matrix}12-(-2) \\ 3-(2.5)\end{matrix}\right)=\left(\begin{matrix}10\\ 5.5\end{matrix}\right)\]

OpenStudy (anonymous):

I forgot to correct ... 12+(-2) and 3+(2.5) => 10 & 5.5